OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of on the given interval. f(t)= t radical 4-t^2 [-1,2]

OpenStudy (anonymous):

ah now we have a question

OpenStudy (anonymous):

take the derivative, check the critical points and the endpoints.

OpenStudy (anonymous):

im having trouble wit the radical

OpenStudy (anonymous):

derivative is found using product rule

OpenStudy (anonymous):

i got to do product rule and chain rule right

OpenStudy (anonymous):

you get \[f'(t)=\sqrt{4-t^2}+\frac{t}{\sqrt{4-t^2}}\times -2t\]

OpenStudy (anonymous):

right product plus chain. then add

OpenStudy (anonymous):

okay i only got up to the derivative but didnt no how to combine

OpenStudy (anonymous):

you get \[\frac{2(2-t^2)}{\sqrt{4-t^2}}\]

OpenStudy (anonymous):

and then just use the numerator equal to zero?

OpenStudy (anonymous):

oh you have to multiply first one top and bottom by \[\sqrt{4-t^2}\]

OpenStudy (anonymous):

so you get \[\frac{4-t^2}{\sqrt{4-t^2}}-\frac{2t^2}{\sqrt{4-t^2}}\]] etc

OpenStudy (anonymous):

okay i get it after that

OpenStudy (anonymous):

wait i messed up sorry

OpenStudy (anonymous):

i forgot the 2 in the denominator of the second one. it should be \[\frac{4-t^2}{\sqrt{4-t^2}}-\frac{t^2}{\sqrt{4-t^2}}\]

OpenStudy (anonymous):

so numerator is \[4-2t^2=2(2-t^2)\] zero at \[\sqrt{2}\]

OpenStudy (anonymous):

also \[-\sqrt{2}\] but that is not in your interval

OpenStudy (anonymous):

okay im good after that

OpenStudy (anonymous):

ok don't forget to evaluate at the endpoints

OpenStudy (anonymous):

just had trouble with the critical point

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw